OpenStudy (anonymous):

I got the equation but I want to double check if it is right or not. Semielliptical Arch Bridge. An arch in the shape of the upper half of an ellipse is used to support a bridge that is to span a river 36 meters wide. The center of the arch is 8 meters above the center of the river. Write an equation for the ellipse in which the x-axis conicides with the water level and then y-axis passes through the center of the arch. Type the left side of the equation of the equation below ____________________= 1

OpenStudy (amistre64):

when x=0; y=8 when y=0; x=36/2

OpenStudy (amistre64):

use those properties to determine the denominators of the standard ellipse equations

OpenStudy (anonymous):

another way of stating what amistre said is: -the major axis is 36 -the minor axis is 16 (8 + 8) -ellipse is centered @ the origin

OpenStudy (anonymous):

should I use the formala x^2/a^2 +y^2/b^2

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ 256 }+\frac{ x ^{2} }{ 64{} }\]

OpenStudy (amistre64):

should you use that formula? yes

OpenStudy (amistre64):

under y is 8^2 and under x is 18^2 so other than a typo, that seems fine

OpenStudy (anonymous):

Thank you!